OpenStudy (anonymous):

In October, Greg and Thomas had the same amount of money in their savings accounts. In November, Greg deposited $120 into his account. Thomas increased the money in his account by 20%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in October? $

OpenStudy (texaschic101):

x = money in savings account in Oct Greg = Thomas x + 120 = x + .20x -- subtract x from both sides x - x + 120 = .20x -- combine like terms 120 = .20x -- divide both sides by .20 120/.20 = x 600 = x check.. Greg -- 600 + 120 = 720 Thomas -- 600 + .20(600) = 600 + 120 = 720 The money they had in the account in Oct. was $ 600 .... and in November, they both had $ 720.